Yogurt blends regular yogurt that is 3% fat with its no fat yogurt to obtain lowfat yogurt that is 1%. How many pounds of regular yogurt and how many pounds of no fat yogurt should be mixed to obtain 60 pounds of low fat yogurt?
Answer by ptaylor(2198)

Let x=amount of regular (3%) yogurt that is needed
Then 60-x=amount of no fat yogurt that is needed
Now we know that the amount of pure fat that exists before the mixture (0.03*x) has to equal the amount of pure fat that exists in the final mixture (0.01*60), so our equation to solve is:
0.03x=0.01*60 simplify
0.03x=0.6 divide each side by 0.03
x=20 lbs-------------------------amount of 3% yogurt needed
60-x=60-20=40 lb-----------amount of no fat yogurt needed
CK
0.03*20+0.0*40=0.1*60
0.6=0.6
Hope this helps---ptaylor